OpenStudy (anonymous):

IS this an acceptable answer? Find the exact value of the expression. 3sin45 degrees - 4tan(pi/6) I said ((3rad2)/2) - (4/rad3) Do I have to combine it or take the radical out of the denominator?

OpenStudy (anonymous):

i would take the radical out of the denominator just to make it look a little nicer.

OpenStudy (anonymous):

what in the world??!!

OpenStudy (anonymous):

right now you have: \[\frac{3\sqrt{2}}{2} - \frac{4}{\sqrt{3}} = \frac{3\sqrt{2}}{2}- \frac{4\sqrt{3}}{3}\] if you wanna go further than that, its your call.

OpenStudy (anonymous):

what kind of idiot math teacher has both degrees and numbers in the same problem?

OpenStudy (anonymous):

lololol, i didnt even notice that.

OpenStudy (anonymous):

let me just shut up. there is no end to bad math instruction

OpenStudy (anonymous):

a teacher that covers his/her bases, im for deg/rad equality!

OpenStudy (anonymous):

like saying square 3 degrees celsius, add 2 degrees kelvin and then subtract the square root of 64 degrees fahrenheit

OpenStudy (anonymous):

hahaha! take it up with my teacher

OpenStudy (anonymous):

its like all the people on this site trying to solve physics problems using g = 9.81 m/s^2 when distances are given in ft >.>

OpenStudy (anonymous):

no just pass the class and hope your next instructor knows some mathematics

OpenStudy (anonymous):

It's just busy work

OpenStudy (anonymous):

this is kinda my next instructor, it's for a summer assignment. so i still have her for a whole year.

OpenStudy (anonymous):

yeah don't take it to heart. @joemath i like the ones where the ball gets hit by a bat and you are supposed to assume it goes straight up!

OpenStudy (anonymous):

@elizac when you feel more confident just ask in a meek tone how come one input is given by numbers and the other with degrees attached? say "just curious but..."

OpenStudy (anonymous):

you should bring it up in class. not in a rude way, be chill about it lol. @Satellite lol, we have some skilled batters haha

OpenStudy (anonymous):

"i just happened to notice that..."

OpenStudy (anonymous):

lol will do!!